### v1

# Approximation of multipartite quantum states: revised version with new applications

Abstract
Special approximation technique for analysis of different characteristics of states of multipartite infinite-dimensional quantum systems is proposed and applied to the study of the relative entropy of $\pi$-entanglement and its regularisation.
In particular, by using this technique we obtain simple sufficient conditions for local continuity (convergence) of the regularized relative entropy of $\pi$-entanglement.
We establish a finite-dimensional approximation property for the relative entropy of entanglement and its regularization that allows us to generalize to the infinite-dimensional case the results proved in the finite-dimensional settings.
We also show that for any multipartite state with finite energy the infimum in the definition of the relative entropy of $\pi$-entanglement can be taken over the set of finitely-decomposable $\pi$-separable states with finite energy.
1 Introduction
A specific feature of infinite-dimensional quantum systems consists in the singular properties (discontinuity, infinite values, etc.) of basic characteristics of quantum states [5, 9, 38]. So, for the strict mathematical analysis of infinite-dimensional quantum systems it is necessary to apply special techniques, in particular, approximation techniques to overcome the problems arising from these singularities.
In this article we develop the approximation technique proposed in [32] in the one-party case to analysis of characteristics of multipartite infinite-dimensional quantum systems. A central notion of this technique is the finite-dimensional approximation property of a state (briefly called the FA-property). According to [32] a quantum state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}$ has the FA-property if there exists a sequence $\{g_{i}\}$ of nonnegative numbers such that
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}g_{i}<+\infty\quad\textrm{and}\quad\lim_{\beta\to 0^{+}}\left[\sum_{i=1}^{+\infty}e^{-\beta g_{i}}\right]^{\beta}=1.$
The FA-property of a state $\rho$ implies the finiteness of its entropy [32], but the converse implication is not true (this is shown in [2] by constructing a subtle example of a state with finite entropy not possessing the FA-property). In Section 2.2 the following simple sufficient condition for the FA-property is proved: the FA-property holds for a state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}$ provided that [^2] This condition strengthens the condition obtained in [32] which states that the FA-property holds for a state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}$ provided that $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{q}i<+\infty$ for some $q>2$.
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}i<+\infty.$
This condition shows that the FA-property holds for all states whose eigenvalues tend to zero faster than $[i\ln^{q}i]^{-1}$ as $i\to+\infty$ for some $q>3$, in particular, it holds for all Gaussian states playing essential role in quantum information theory.
Many characteristics of a state of a $n$-partite quantum system $A_{1}...A_{n}$ have the form of a function
$f(\rho\,|\,p_{1},...,p_{l})$
on the set $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ of states of this system depending on some parameters $p_{1},...,p_{l}$ (other states, quantum channels, quantum measurements, etc.). If $\rho$ is a state of infinite-dimensional $n$-partite quantum system $A_{1}...A_{n}$ then we may approximate it by the sequence of states
$\rho_{r}=Q_{r}\rho\hskip 1.0ptQ_{r}[\mathrm{Tr}Q_{r}\rho\hskip 1.0pt]^{-1},\quad Q_{r}=P_{r}^{1}\otimes...\otimes P_{r}^{n},$ (1)
where $P_{r}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account). Naturally, the question arises under what conditions
$f(\rho_{r}\,|\,p_{1},...,p_{l})\quad\textrm{tends to}\quad f(\rho\,|\,p_{1},...,p_{l})\quad\textrm{as}\;\;r\to\infty$
uniformly on $p_{1},...,p_{l}$. Our main technical result asserts that for given $m\leq n$ the FA-property of the marginal states $\rho_{A_{1}}$,…, $\rho_{A_{m}}$ of a state $\rho$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ guarantees this uniform convergence for a wide class of functions $f$ (depending on $m$). In fact, the more general assertion is valid in which the states $\rho_{r}$ are defined by formula (1) via the operators $Q_{r}=P_{r}^{s_{1}}\otimes...\otimes P_{r}^{s_{l}}\otimes I_{R}$, where $\{s_{1},...,s_{l}\}$ is any subset of $\{1,...,n\}$ and $R=A_{1}...A_{n}\setminus A_{s_{1}}...A_{s_{l}}$. For a smaller class of functions on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ the above result is valid with the FA-property of the states $\rho_{A_{1}}$,…, $\rho_{A_{m}}$ replaced by the weaker condition of finiteness of their entropy (Theorem 3 in Section 3).
[…]
The proposed approximation technique allows us to establish the lower semicontinuity of the regularized relative entropy of $\pi$-entanglement $E^{\infty}_{R,\pi}$ on the subset $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ of $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of states $\rho$ that have at least $n-1$ marginal states with the FA-property. Moreover, it is proved that
$\liminf_{k\to+\infty}E^{\infty}_{R,\pi}(\rho_{k})\geq E^{\infty}_{R,\pi}(\rho_{0}),$
for arbitrary sequence $\{\rho_{k}\}$ converging to a state $\rho_{0}$ in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$.
By combining this observation with the results from [15] we obtain simple sufficient conditions for local continuity of the regularized relative entropy of $\pi$-entanglement $E^{\infty}_{R,\pi}$, i.e. conditions under which the limit relation
$\lim_{k\to+\infty}E^{\infty}_{R,\pi}(\rho_{k})=E^{\infty}_{R,\pi}(\rho_{0})<+\infty$
holds for a given sequence $\{\rho_{k}\}$ converging to a state $\rho_{0}$.
The approximation technique is used also to obtain several observations concerning definition of the relative entropy of $\pi$-entanglement. In particular, we prove that for all states in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ this quantity can be defined as the relative entropy distance to the set of all finitely decomposable separable states (despite the existence of finitely-non-decomposable and countably-non-decomposable separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$).
We establish a finite-dimensional approximation property for the relative entropy of entanglement and its regularization in $n$-partite quantum systems that allows us to generalize to the infinite-dimensional case the results proved in the finite-dimensional settings. Examples of using this property are presented.
Finally, we consider energy-constrained versions of the $n$-partite relative entropy of $\pi$-entanglement. It is proved, in particular, that for any state in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ with finite energy the infimum in the definition of the relative entropy of $\pi$-entanglement can be taken over all finitely-decomposable separable states with finite energy (provided that the Hamiltonians of individual subsystems satisfy a certain condition).
2 Preliminaries
2.1 Basic notation
[…]
Write $I_{\mathcal{H}}$ for the unit operator on a Hilbert space $\mathcal{H}$ and $\mathrm{Id}_{\mathcal{\mathcal{H}}}$ for the identity transformation of the Banach space $\mathfrak{T}(\mathcal{H})$.
The von Neumann entropy of a quantum state $\rho\in\mathfrak{S}(\mathcal{H})$ is defined by the formula $S(\rho)=\operatorname{Tr}\eta(\rho)$, where $\eta(x)=-x\ln x$ for $x>0$ and $\eta(0)=0$. It is a concave lower semicontinuous function on the set $\mathfrak{S}(\mathcal{H})$ taking values in $[0,+\infty]$ [9, 16, 38]. The von Neumann entropy satisfies the inequality
$S(p\rho+(1-p)\sigma)\leq pS(\rho)+(1-p)S(\sigma)+h_{2}(p)$ (2)
valid for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H})$ and $p\in(0,1)$ with possible value +$\infty$+ ∞ in both sides, where $\,h_{2}(p)=\eta(p)+\eta(1-p)\,$ is the binary entropy [18, 39].
The quantum relative entropy for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H})$ can be defined as
$D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma)=\sum_{i}\langle\varphi_{i}|\,\rho\ln\rho-\rho\ln\sigma\,|\varphi_{i}\rangle,$
where $\{\varphi_{i}\}$ is the orthonormal basis of eigenvectors of the state $\rho$ and it is assumed that $D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma)=+\infty$ if $\,\mathrm{supp}\rho\hskip 1.0pt$ is not contained in $\hskip 1.0pt\mathrm{supp}\hskip 1.0pt\sigma$ [16, 9, 39]. [^3] The support $\mathrm{supp}\rho$ of a positive trace class operator $\rho$ is the closed subspace spanned by the eigenvectors of $\rho$ corresponding to its positive eigenvalues.
The quantum conditional entropy
$S(A|B)_{\rho}=S(\rho)-S(\rho_{\hskip 1.0ptB})$
of a state $\rho$ of a bipartite quantum system $AB$ with finite marginal entropies $S(\rho_{A})$ and $S(\rho_{B})$ is essentially used in analysis of quantum systems [9, 39]. It can be extended to the set of all states $\rho$ with finite $S(\rho_{A})$ by the formula
$S(A|B)_{\rho}=S(\rho_{A})-D(\rho\hskip 1.0pt\|\hskip 1.0pt\rho_{A}\otimes\rho_{B})$ (3)
proposed in [12], where it is shown that this extension possesses all basic properties of the quantum conditional entropy valid in finite dimensions.
The quantum mutual information (QMI) of a state $\,\rho\,$ of a multipartite quantum system $A_{1}\ldots A_{n}$ is defined as (cf.[17, 8, 40])
$I(A_{1}\!:\ldots:\!A_{n})_{\rho}\doteq D(\rho\,\|\,\rho_{A_{1}}\otimes\cdots\otimes\rho_{A_{n}})=\sum_{s=1}^{n}S(\rho_{A_{s}})-S(\rho),$
where the second formula is valid if $S(\rho)<+\infty$. It is easy to show (see, f.i., [31]) that
$I(A_{1}\!:\ldots:\!A_{n})_{\rho}\leq 2\sum_{s=1}^{n-1}S(\rho_{A_{s}}).$ (4)
Similar upper bound holds with any other $n-1$ marginal entropies of the state $\rho$.
For any positive (semi-definite) densely defined operator [^4] We assume that a positive operator is a self-adjoint operator [24]. $H$ on a Hilbert space $\mathcal{H}$ and any positive operator $\rho$ in $\mathfrak{T}(\mathcal{H})$ we assume that
$\mathrm{Tr}H\rho=\sup_{n}\mathrm{Tr}P_{n}H\rho\leq+\infty,$ (5)
where $P_{n}$ is the spectral projector of $H$ corresponding to the interval $[0,n]$. Then
$\mathfrak{C}_{H,E}=\left\{\rho\in\mathfrak{S}(\mathcal{H})\,|\,\mathrm{Tr}H\rho\leq E\right\}$ (6)
is a closed convex nonempty subset of $\mathfrak{S}(\mathcal{H})$ for any $E$ exceeding the infimum of the spectrum of $H$. If $H$ is treated as Hamiltonian of a quantum system associated with the space $\mathcal{H}$ then $\mathfrak{C}_{H,E}$ is the set of states with the mean energy not exceeding $E$.
We will pay a special attention to the class of unbounded densely defined positive operators on $\mathcal{H}$ having discrete spectrum. In Dirac’s notations any such operator $H$ can be represented as follows
$H=\sum_{i=1}^{+\infty}h_{i}|\tau_{i}\rangle\langle\tau_{i}|$ (7)
on the domain $\mathcal{D}(H)=\{\varphi\in\mathcal{H}\,|\,\sum_{i=1}^{+\infty}h^{2}_{i}|\langle\tau_{i}|\varphi\rangle|^{2}<+\infty\}$, where $\left\{\tau_{i}\right\}_{i=1}^{+\infty}$ is the orthonormal basis of eigenvectors of $H$ corresponding to the nondecreasing sequence $\left\{\smash{h_{i}}\right\}_{i=1}^{+\infty}$ of eigenvalues tending to +$\infty$+ ∞.
It is well known that the von Neumann entropy is continuous on the set $\mathfrak{C}_{H,E}$ for any $E$ if (and only if) the operator $H$ satisfies the condition
$\mathrm{Tr}\,e^{-\beta H}<+\infty\quad\textrm{for all}\;\beta>0$ (8)
and that the maximal value of the entropy on this set is achieved at the Gibbs state $\gamma_{H}(E)\doteq e^{-\beta(E)H}/\mathrm{Tr}e^{-\beta(E)H}$, where the parameter $\beta(E)$ is determined by the equality $\mathrm{Tr}He^{-\beta(E)H}=E\mathrm{Tr}e^{-\beta(E)H}$ [38]. Condition (8) implies that $H$ is an unbounded operator having discrete spectrum, i.e. it has the form (7). So, by the Lemma in [10] the set $\mathfrak{C}_{H,E}$ defined in (6) is compact for any $E$.
We will use the function $F_{H}(E)\doteq S(\gamma_{H}(E))=\max_{\rho\in\mathfrak{C}_{H,E}}S(\rho)$. By Proposition 1 in [26] condition (8) holds if and only if
$F_{H}(E)=o\hskip 1.0pt(E)\quad\textrm{as}\quad E\rightarrow+\infty.$ (9)
We will use the following
Proposition 1.  Let $\rho$ be a state in $\mathfrak{S}(\mathcal{H})$. The following properties are equivalent:
  1. (i)
    $S(\rho)<+\infty$;
  2. (ii)
    there is a positive operator $H$ on $\mathcal{H}$ satisfying condition (8) s.t. $\mathrm{Tr}H\rho<+\infty$;
  3. (iii)
    $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln i<+\infty$, where $\{\lambda^{\rho}_{i}\}_{i=1}^{+\infty}$ is the sequence of eigenvalues of $\rho$ arranged in the non-increasing order.
If equivalent properties $(i)\textrm{-}(iii)$ hold then the positive operator $H$ in $(ii)$ can be taken diagonizable in any basis of eigenvectors of the state $\rho$.
Proof. The implication $\rm(ii)\Rightarrow(i)$ follows from the results mentioned before. To prove $\rm(i)\Rightarrow(iii)$ it suffices to note that $\lambda^{\rho}_{i}\leq 1/i$ for all $i$.
Assume that $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln i<+\infty$. Then it is easy to show the existence of an increasing sequence $\{c_{i}\}_{i=1}^{+\infty}$ such that $\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}c_{i}\ln i<+\infty$. Then $\rm(ii)$ holds with the operator
$H=\sum_{i=1}^{+\infty}c_{i}\ln i|\varphi_{i}\rangle\langle\varphi_{i}|,$
where $\{\varphi_{i}\}_{i=1}^{+\infty}$ is any basis of eigenvectors of $\rho$ corresponding to the sequence $\{\lambda^{\rho}_{i}\}_{i=1}^{+\infty}$. $\Box$
We will often consider positive operators $H$ satisfying the condition
$\lim_{\beta\rightarrow 0^{+}}\left[\mathrm{Tr}\,e^{-\beta H}\right]^{\beta}=1,$ (10)
which is stronger than condition (8). The condition (10) is valid if $\;\liminf\limits_{i\rightarrow+\infty}h_{i}/\ln^{p}i>0\,$ for some $\,p>2$, where $\{h_{i}\}$ is the sequence of eigenvalues of $H$ [28, Proposition 1]. By Lemma 1 in [28] condition (10) holds if and only if
$F_{H}(E)=o\hskip 1.0pt(\sqrt{E})\quad\textrm{as}\quad E\rightarrow+\infty$ (11)
(this should be compared with (9)). It is essential that condition (10) is valid for the Hamiltonians of many real quantum systems [1, 28].
We will use the following
Lemma 1. If $H$ is a positive operator on $\mathcal{H}$ satisfying condition (8) then $H^{2}$ is a positive operator on $\mathcal{H}$ satisfying condition (10).
[…]
$F_{H^{2}}(E)\leq F_{H}(\sqrt{E})\quad\forall E\geq E^{2}_{0},$
where $E_{0}$ is the minimal eigenvalue of $H$. This inequality follows from the definitions of $F_{H^{2}}(E)$ and $F_{H}(\sqrt{E})$ due to the inequality
$\mathrm{Tr}H\rho\leq\sqrt{\mathrm{Tr}H^{2}\rho}$
valid for any state $\rho$ in $\mathfrak{S}(\mathcal{H})$ (with possible value +$\infty$+ ∞ in one or both sides). The last inequality is a direct corollary of the concavity of the square root, since
$\mathrm{Tr}H\rho=\sum_{i}h_{i}p_{i}\quad\textrm{and}\quad\mathrm{Tr}H^{2}\rho=\sum_{i}h^{2}_{i}p_{i},\quad p_{i}=\langle\tau_{i}|\rho|\tau_{i}\rangle,$
where it is assumed that the operator $H$ has the form (7). $\Box$
2.2 Simple sufficient condition for the FA-property
Following [32] we will say that a quantum state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}_{i=1}^{+\infty}$ has the FA-property if there exists a sequence $\{g_{i}\}_{i=1}^{+\infty}$ of nonnegative numbers such that
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}g_{i}<+\infty\quad\textrm{and}\quad\lim_{\beta\to 0^{+}}\left[\sum_{i=1}^{+\infty}e^{-\beta g_{i}}\right]^{\beta}=1.$ (12)
We will write $\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H})$ for the set of all states in $\mathfrak{S}(\mathcal{H})$ having the FA-property. It follows from Corollary 3 in [32] that $\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H})$ is a face [^5] It means that the set $\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H})$ is convex and contains any segment from $\mathfrak{S}(\mathcal{H})$ provided that it contains at least one internal point of this segment. of the convex set $\mathfrak{S}(\mathcal{H})$ contained in the face of all states in $\mathfrak{S}(\mathcal{H})$ with finite entropy (these two faces do not coincide by the examples in [2]).
The following proposition gives a simple sufficient condition for the FA-property.
Proposition 2.  A quantum state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}_{i=1}^{+\infty}$ has the FA-property provided that
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}i<+\infty.$ (13)
Proof. Let $\rho=\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}|\varphi_{i}\rangle\langle\varphi_{i}|$ be a spectral decomposition of $\rho$.
If (13) holds then there exists an increasing sequence $\{c_{i}\}_{i=1}^{+\infty}$ of positive numbers tending to +$\infty$+ ∞ such that
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}c_{i}\ln^{2}i<+\infty.$ (14)
It is easy to see that the operator
$H=\sum_{i=1}^{+\infty}\sqrt{c_{i}}\ln i|\varphi_{i}\rangle\langle\varphi_{i}|$
satisfies condition (8). By Lemma 2.1 the operator $H^{2}$ satisfies condition (10). Since (14) implies that $\mathrm{Tr}H^{2}\rho<+\infty$, the state $\rho$ has the FA-property by the Theorem in [32, Section 2]. $\Box$
Remark 1. The condition given by Proposition 2.2 is strictly stronger than the last claim of the Theorem in [32] which states that a quantum state $\rho$ with the spectrum $\{\lambda^{\rho}_{i}\}_{i=1}^{+\infty}$ has the FA-property provided that
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{q}i<+\infty\quad\textrm{for some}\quad q>2.$ (15)
Indeed, it is easy to see that any state with the spectrum $\{c[i\ln i\ln^{p}(\ln i)]^{-1}\}_{i=10}^{+\infty}$, $p>1$, where $c$ is a normalizing constant, satisfies condition (13) but does not satisfy condition (15).
Remark 2. If $\rho$ is a state in $\mathfrak{S}(\mathcal{H})$ with the spectrum $\{\lambda^{\rho}_{i}\}_{i=1}^{+\infty}$ arranged in the non-increasing order then Proposition 2.2 show that
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}\lambda^{\rho}_{i}<+\infty\;\Rightarrow\;\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}i<+\infty\;\Rightarrow\;\rho\in\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H})\;\Rightarrow\;S(\rho)<+\infty,$ (16)
where $\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H})$ denotes the subset of states having the FA-property (the first implication in (16) follows from the inequality $\lambda^{\rho}_{i}\leq 1/i,\;\forall i$, the third one is mentioned in the Theorem in [32, Section 2]). In [2] it is shown that the third implication in (16) cannot be replaced by """"$\Leftrightarrow$"" ⇔ " by constructing examples of states with finite entropy which do not have the FA-property. We may conjecture (until proven otherwise) that the second implication in (16) can be replaced by """"$\Leftrightarrow$"" ⇔ ", i.e. that (13) is a criterion of the FA-property of a state with the spectrum $\{\lambda^{\rho}_{i}\}$ arranged in the non-increasing order.
2.3 On uniform continuity under the energy-type constraints
[…]
Assume that $H_{A_{1}}$,…,$H_{A_{m}}$ are arbitrary positive densely defined operators on the spaces $\mathcal{H}_{A_{1}}$,…,$\mathcal{H}_{A_{m}}$, $m\leq n$. For a given function $f$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, $\varepsilon\in(0,1]$ and $E>0$ introduce the quantity
$B_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})\doteq\sup\left\{|f(\rho)-f(\sigma)|\,\left|\,\rho,\sigma\in\mathfrak{C}(E|H_{A_{1}},...,H_{A_{m}}),\textstyle\frac{1}{2}\|\rho-\sigma\|_{1}\leq\varepsilon\right.\right\}\!,$
[…]
$\mathfrak{C}(E|H_{A_{1}},...,H_{A_{m}})\doteq\left\{\varrho\in\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})\,\left|\,\sum_{k=1}^{m}\mathrm{Tr}H_{A_{k}}\varrho_{A_{k}}\leq mE\right.\right\}.$ (17)
The quantity $B_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})$ is the (optimal) modulus of continuity of the function $f$ under the energy-type constraint induced by the operators $H_{A_{1}}$,…,$H_{A_{m}}$.
Definition 1.  A function $f$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ belongs to the class $I_{n}^{m}$ if
$\lim_{\varepsilon\to 0}B_{f}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})=0$
for any $E>0$ and arbitrary positive operators $H_{A_{1}}$,….,$H_{A_{m}}$ satisfying condition (8).
Definition 2.  A function $f$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ belongs to the class $II_{n}^{m}$ if
$\lim_{\varepsilon\to 0}B_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})=0$
for any $E>0$ and arbitrary positive operators $H_{A_{1}}$,…,$H_{A_{m}}$ satisfying condition (10).
According to these definitions the class $I_{n}^{m}$ (resp. the class $II_{n}^{m}$) consists of all functions uniformly continuous on the set $\mathfrak{C}(E|H_{A_{1}},...,H_{A_{m}})$ (defined in (17)) for any $E>0$ and arbitrary positive operators $H_{A_{1}}$,…,$H_{A_{m}}$ satisfying condition (8) (resp. condition (10)). Since condition (10) is stronger than condition (8), we have $I_{n}^{m}\subseteq II_{n}^{m}$.
Let $f$ be a given function from the class $I_{n}^{m}$ (resp. $II_{n}^{m}$). We will say that $\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})$ is a faithful $I_{n}^{m}$-type (resp. $II_{n}^{m}$-type) continuity bound for the function $f$ provided that
$B_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})\leq\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})\quad\forall E>E_{0},\,\varepsilon\in(0,1],$
and
$\lim_{\varepsilon\to 0}\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})=0\quad\forall E>E_{0}$
for arbitrary positive operators $H_{A_{1}}$,…,$H_{A_{m}}$ satisfying condition (8) (resp. condition (10)), where $mE_{0}$ is the sum of the minimal eigenvalues of the operators $H_{A_{1}}$,…,$H_{A_{m}}$.
Let $\mathcal{F}$ be a family of functions from the class $I_{n}^{m}$ (resp. $II_{n}^{m}$). We will say that $\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})$ is a common faithful $I_{n}^{m}$-type (resp. $II_{n}^{m}$-type) continuity bound for the family $\mathcal{F}$ provided that
$B_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})\leq\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})\quad\forall f\in\mathcal{F},\;E>E_{0},\;\varepsilon\in(0,1]$
and
$\lim_{\varepsilon\to 0}\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})=0\quad\forall E>E_{0}$
for arbitrary positive operators $H_{A_{1}}$,…,$H_{A_{m}}$ satisfying condition (8) (resp. condition (10)).
Among characteristics of a composite infinite-dimensional quantum system $A_{1}...A_{n}$ there are many that belong to the classes $I_{n}^{m}$ and $II_{n}^{m}$ for some $m\leq n$ [^7] It follows from the definitions that if $f\in I_{n}^{m}$ (resp. $f\in II_{n}^{m}$) for some $m<n$ then $f\in I_{n}^{u}$ (resp. $f\in II_{n}^{u}$) for any $u=\overline{m,n}$. For example, the von Neumann entropy of a state of $A_{1}$ and the quantum conditional entropy $S(A_{1}|A_{2})$ belong, respectively, to the classes $I_{1}^{1}$ and $I^{1}_{2}$. The $I_{1}^{1}$-type and $I^{1}_{2}$-type continuity bounds for these characteristics are obtained by Winter in [41]. The $I^{1}_{2}$-type continuity bounds for the quantum mutual information $I(A_{1}\!:\!A_{2})$, for the entanglement of formation $E_{F}$ and for the squashed entanglement $E_{sq}$ in a bipartite system $A_{1}A_{2}$ obtained in [27] by using Winter’s technique show that these characteristics belong to the class $I^{1}_{2}$. Moreover, the last two entanglement measures have a common faithful $I^{1}_{2}$-type continuity bound. The regularized entanglement of formation $E^{\infty}_{F}$ also belongs to the class $I^{1}_{2}$: the $I^{1}_{2}$-type continuity bounds for $E^{\infty}_{F}$ obtained by using Winter’s technique is presented in Proposition 18 in [33]. [^8] We do not distinguish between the discrete and continuous versions of $E_{F}$ and $E^{\infty}_{F}$, since these versions coincide at any state $\omega\in\mathfrak{S}(\mathcal{H}_{A_{1}A_{2}})$ such that at least one of the states $\omega_{A_{1}}$ and $\omega_{A_{2}}$ has finite energy w.r.t. a Hamiltonian satisfying condition (8) [33, Section 4.4.2].
The universal version of the Alicki-Fannes-Winter technique presented in [33] allows us to show that
$\bigcup_{C,D\geq 0}\left[\widehat{L}^{m}_{n}(C,D)\cup N^{m}_{n,1}(C,D)\cup N^{m}_{n,2}(C,D)\cup N^{m}_{n,3}(C,D)\right]\subset II_{n}^{m},$ (18)
where $\widehat{L}^{m}_{n}(C,D)$ and $N^{m}_{n,s}(C,D)$, $s=1,2,3$, are the classes of functions on $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ introduced in [31, 33] which contain many important characteristics of composite quantum systems. [^9] A non-complete list of these characteristics and detailed description of the classes $\widehat{L}^{m}_{n}(C,D)$ and $N^{m}_{n,s}(C,D)$, $s=1,2,3$, can be found in the Appendix. Moreover, Theorem 6 in [33] implies that all functions in the family
$\mathcal{F}(C,D)=\widehat{L}^{m}_{n}(C,D)\cup N^{m}_{n,1}(C,D)\cup N^{m}_{n,2}(C,D)\cup N^{m}_{n,3}(C,D),\quad C,D\geq 0,$
have the common faithful $II_{n}^{m}$-type continuity bound
$\!\widehat{B}_{\mathcal{F}(C,D)}(\varepsilon,E|H_{A_{1}},..,H_{A_{m}})=C\sqrt{\varepsilon(2-\varepsilon)}F_{H_{A_{1}},..,H_{A_{m}}}\!\!\left[\frac{2mE}{\varepsilon(2-\varepsilon)}\right]+Dg(\sqrt{\varepsilon(2-\varepsilon)}),$ (19)
where
$F_{H_{A_{1}},...,H_{A_{m}}}(E)=\sup\left\{S(\rho)\,\left|\,\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}..A_{m}}),\,\sum_{k=1}^{m}\mathrm{Tr}H_{A_{k}}\rho_{A_{k}}\leq E\right.\right\}$ (20)
and
$g(x)=(x+1)h_{2}\!\left(\frac{x}{x+1}\right)=(x+1)\ln(x+1)-x\ln x,\;\,x>0,\quad g(0)=0.$ (21)
The faithfulness of the common continuity bound (19) for the family $\mathcal{F}(C,D)\subset II_{n}^{m}$ follows from the fact that the validity of condition (10) for all the operators $H_{A_{1}},...,H_{A_{m}}$ implies that
$F_{H_{A_{1}},...,H_{A_{m}}}(E)=o(\sqrt{E})\quad\textrm{as}\quad E\to+\infty$
by Lemma 4 in [31].
We will see in the next Section 3 that belonging of a given characteristic to the class $II_{n}^{m}$ allows us to realise effective finite-dimensional approximation of this characteristic. If, moreover, this characteristic belongs to the class $I_{n}^{m}$ then we may relax some requirements in such approximation (briefly speaking, to replace the FA-property of marginal states by the finiteness of the von Neumann entropy of these states).
Practically, to show that a given characteristic $f$ belongs to the class $II_{n}^{m}$ we have either to use general relation (18) or to construct an explicit continuity bound $\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ for $f$ which tends to zero as $\varepsilon\to 0$ for arbitrary positive operators $H_{A_{1}}$,….,$H_{A_{m}}$ satisfying condition (10). More efforts are required to show that a given characteristic $f$ belongs to the class $I_{n}^{m}$: it is necessary to construct an explicit continuity bound $\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ for $f$ which tends to zero as $\varepsilon\to 0$ for arbitrary positive operators $H_{A_{1}}$,….,$H_{A_{m}}$ satisfying condition (8). Sometimes one can use the following simple
Lemma 2. If $f$ and $g$ are functions from the class $I_{n}^{m}$ (resp. $II_{n}^{m}$) then any their linear combitation is a function from the class $I_{n}^{m}$ (resp. $II_{n}^{m}$).
Example 1. The one-way classical correlation $C_{B}$ and the quantum discord $D_{B}$ in a bipartite infinite-dimensional system $AB$ with measured system $B$ (cf.[7, 20, 35]) belong to class $II_{2}^{1}$ in both settings $A_{1}=A,A_{2}=B$ and $A_{1}=B,A_{2}=A$. This follows from (18), since $C_{B}\in\widehat{L}_{2}^{1}(1,2)$ in both settings $A_{1}=A,A_{2}=B$ and $A_{1}=B,A_{2}=A$ and $D_{B}\in\widehat{L}_{2}^{1}(C,2)$, where $C=2$ in the setting $A_{1}=A,A_{2}=B$ and $C=1$ in the setting $A_{1}=B,A_{2}=A$ [33, Section 4.3]. The $II_{2}^{1}$-type continuity bounds for the regularized one-way classical correlation $C^{\infty}_{B}$ presented in [33, Proposition 16] implies that $C^{\infty}_{B}\in II_{2}^{1}$ in both settings $A_{1}=A,A_{2}=B$ and $A_{1}=B,A_{2}=A$.
More detailed analysis shows that the functions $C_{B}$, $C^{\infty}_{B}$ and $D_{B}$ belong to class $I_{2}^{1}$ in the setting $A_{1}=A,A_{2}=B$. Indeed, by the Koashi-Winter relations (cf.[13]) we have
$C_{B}(\omega_{AB})=S(\omega_{A})-E_{F}(\omega_{AC})\quad\textrm{and}\quad C_{B}^{\infty}(\omega_{AB})=S(\omega_{A})-E^{\infty}_{F}(\omega_{AC})$ (22)
for any pure state $\omega$ in $\mathfrak{S}(\mathcal{H}_{ABC})$ such that $S(\omega_{A})<+\infty$, where $E_{F}$ and $E_{F}^{\infty}$ are the entanglement of formation and its regularization.
By applying the faithful $I_{1}^{1}$-type continuity bound for the entropy and the $I_{2}^{1}$-type continuity bounds for the entanglement of formation and its regularization mentioned before (18) it is easy to show that the functions $\omega\mapsto S(\omega_{A})$, $\omega\mapsto E_{F}(\omega_{AC})$ and $\omega\mapsto E^{\infty}_{F}(\omega_{AC})$ on $\mathfrak{S}(\mathcal{H}_{ABC})$ have faithful $I_{3}^{1}$-type continuity bounds in the setting $A_{1}=A,A_{2}=B,A_{3}=C$.
Since for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H}_{AB})$ such that $\frac{1}{2}\|\rho-\sigma\|_{1}\leq\varepsilon$ there exist pure states $\hat{\rho}$ and $\hat{\sigma}$ in $\mathfrak{S}(\mathcal{H}_{ABC})$ such that $\frac{1}{2}\|\hat{\rho}-\hat{\sigma}\|_{1}\leq\sqrt{\varepsilon(2-\varepsilon)}$ [9, 39], using the continuity bounds mentined before, the expressions in (22) and Lemma 2.3 one can prove that $C_{B}$ and $C^{\infty}_{B}$ belong to class $I_{2}^{1}$ in the setting $A_{1}=A,A_{2}=B$.
To prove that $D_{B}$ belongs to class $I_{2}^{1}$ in the setting $A_{1}=A,A_{2}=B$, it suffice to note that $D_{B}=I(A\!:\!B)-C_{B}$ and that $I(A\!:\!B)$ belongs to class $I_{2}^{1}$ in the same setting (as mentioned before) and to use Lemma 2.3.
The question of whether the functions $C_{B}$, $C^{\infty}_{B}$ and $D_{B}$ belong to class $I_{2}^{1}$ in the setting $A_{1}=B,A_{2}=A$ remains open. $\Box$
[…]
$I(A_{1}\!:\!...\!:\!A_{n})_{\omega}=I(A_{n-1}\!:\!A_{n})_{\omega}+I(A_{n-2}\!:\!A_{n-1}A_{n})_{\omega}+...+I(A_{1}\!:\!A_{2}...A_{n})_{\omega}$ (23)
valid for any state $\omega\in\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ and Lemma 2.3 it is easy to show that the $n$-partite quantum mutual information $I(A_{1}\!:\!...\!:\!A_{n})$ belongs to class $I_{n}^{n-1}$.
By the same way one can show that the (extended) quantum conditional mutual information $I(A_{1}\!:\!...\!:\!A_{n}|A_{n+1})$ (defined by the rule described in [33, Section 4.2]) belongs to class $I_{n+1}^{n-1}$ (since $I(A_{1}\!:\!A_{2}|A_{3})$ lies in $I_{3}^{1}$ due to the $I_{3}^{1}$-type continuity bound for this characteristic obtained in [27]). $\Box$
[…]
3 Approximation of multipartite quantum states
In this section we obtain our main result about approximation of characteristics of $n$-partite quantum systems belonging to the classes $I_{n}^{m}$ and $II_{n}^{m}$ described in Section 2.3 (Definitions 2.3 and 2.3).
Introduce the approximation map $\Lambda^{s_{1},...,s_{l}}_{r}$ on the space $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ determined by a given set $\{s_{1},...,s_{l}\}\subseteq\{1,...,n\}$ as follows
$\Lambda^{s_{1},...,s_{l}}_{r}(\rho)=Q_{r}\rho\hskip 1.0ptQ_{r}[\mathrm{Tr}Q_{r}\rho\hskip 1.0pt]^{-1},\quad Q_{r}=P_{r}^{s_{1}}\otimes...\otimes P_{r}^{s_{l}}\otimes I_{R},$ (24)
where $P_{r}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account) and $R=A_{1}...A_{n}\setminus A_{s_{1}}...A_{s_{l}}$.
Theorem 1.  Let $A_{1}$,.,$A_{n}$ be arbitrary infinite-dimensional quantum systems, $n\geq 1$.
A) Let $\rho$ be a state in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ such that $S(\rho_{A_{s}})<+\infty$ for $s=\overline{1,m}$, $m\leq n$. Let $\mathcal{F}\subset I_{n}^{m}$ be a family of functions on $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ having common faithful $I^{m}_{n}$-type continuity bound. [^11] See Section 2.3 (Definition 2.3 and below). Then for a given subset $\{s_{1},...,s_{l}\}$ of $\{1,...,n\}$ there is a vanishing sequence $\{Y(r)\}_{r=1}^{+\infty}$ such that
$|f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))-f(\rho)|\leq Y(r)$ (25)
for all $r\in\mathbb{N}$ and all functions $f\in\mathcal{F}$.
[…]
$|f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))-f(\rho)|\leq Y(r)$ (26)
for all $r\in\mathbb{N}$ and all functions $f\in\mathcal{F}$.
The sequence $\{Y(r)\}_{r=1}^{+\infty}$ in claims A and B is completely determined by the states $\rho_{A_{s}}$, $s\in\{s_{1},...,s_{l}\}\cup\{1,...,m\}$, and the common continuity bound for the family $\mathcal{F}$.
Remark 3. In many applications of Theorem 3 the family $\mathcal{F}$ consists of a single function $f$. In this case it suffices to require that $f\in I_{n}^{m}$ in claim A and $f\in II_{n}^{m}$ in claim B.
Remark 4. Claim B of Theorem 3 differs from claim A by the weaker assumption about the family $\mathcal{F}$ (the class $II_{n}^{m}$ instead of $I_{n}^{m}$) which is compensated by the stronger assumption about the marginal states $\rho_{A_{s}}$, $s=\overline{1,m}$ (the FA-property instead of the finiteness of the entropy).
Claim B is more applicable to practical tasks, since the list of characteristics for which faithful $I_{n}^{m}$-type continuity bounds have been obtained (so far) is very short (see Section 2.3) while relation (18) (along with the observation in Section V in [31]) allows us to apply claim B to many characteristics of $n$-partite quantum systems.
Remark 5. The map $\Lambda^{s_{1},...,s_{l}}_{r}$ defined in (24) is nonlinear. The arguments from the proof of Theorem 3 show that its claims remain valid if we replace the family of maps $\Lambda^{s_{1},...,s_{l}}_{r}$ with the family of product-channels
$\widetilde{\Lambda}^{s_{1},...,s_{l}}_{r}(\rho)=\Phi^{s_{1}}_{r}\otimes...\otimes\Phi^{s_{l}}_{r}\otimes\mathrm{Id}_{R}(\rho),\quad\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}}),$ (27)
where $\Phi^{s}_{r}(\varrho)=P_{r}^{s}\varrho P_{r}^{s}+[\mathrm{Tr}(I_{A_{s}}-P_{r}^{s})\varrho]\tau_{s}$ is a channel from $\mathfrak{T}(\mathcal{H}_{A_{s}})$ to itself, $P_{r}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account), $\tau_{s}$ is a pure state in $\mathfrak{S}(\mathcal{H}_{A_{s}})$ corresponding to the maximal eigenvalue of $\rho_{A_{s}}$ and $R=A_{1}...A_{n}\setminus A_{s_{1}}...A_{s_{l}}$.
Proof of Theorem 3. A) For each natural $s$ in $[1,m]$ let $\{\varphi^{s}_{i}\}_{i=1}^{+\infty}$ be the basis of eigenvectors of the state $\rho_{A_{s}}$ corresponding to the non-increasing sequence $\{\lambda^{s}_{i}\}_{i=1}^{+\infty}$ of its eigenvalues. By Proposition 2.1 in Section 2.1 there is a positive operator
$G_{\!A_{s}}=\sum_{i=1}^{+\infty}g^{s}_{i}|\varphi^{s}_{i}\rangle\langle\varphi^{s}_{i}|$ (28)
on $\mathcal{H}_{A_{s}}$ satisfying condition (8) such that $g_{1}^{s}=0$ and
$E_{s}\doteq\mathrm{Tr}G_{\!A_{s}}\rho_{A_{s}}=\sum_{i=1}^{+\infty}\lambda^{s}_{i}g^{s}_{i}<+\infty.$
For each natural $s$ in $[1,n]$ let $P_{r}^{s}=\sum_{i=1}^{r}|\varphi^{s}_{i}\rangle\langle\varphi^{s}_{i}|$ be the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account) and $\bar{P}_{r}^{s}=I_{A_{s}}-P_{r}^{s}$. Then
$\mathrm{Tr}Q_{r}\rho\geq 1-\sum_{j=1}^{l}\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}},$ (29)
where $Q_{r}$ is the projector defined in (24). To prove this inequality note that
$\left|\mathrm{Tr}Q_{r}^{j-1}\rho-\mathrm{Tr}Q_{r}^{j}\rho\hskip 1.0pt\right|\leq\|Q_{r}^{j-1}\|\mathrm{Tr}[\bar{P}_{r}^{s_{j}}\otimes I_{A_{1}...A_{n}\setminus A_{s_{j}}}]\rho=\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}},\quad j=\overline{1,l},$
where $Q_{r}^{0}=I_{A_{1}..A_{n}}$, $\,Q_{r}^{j}=P_{r}^{s_{1}}\otimes...\otimes P_{r}^{s_{j}}\otimes I_{R_{j}}$, $R_{j}=A_{1}...A_{n}\setminus A_{s_{1}}...A_{s_{j}}$. Hence
$1-\mathrm{Tr}Q_{r}\rho\leq\sum_{j=1}^{l}\left|\mathrm{Tr}Q_{r}^{j-1}\rho-\mathrm{Tr}Q_{r}^{j}\rho\hskip 1.0pt\right|\leq\sum_{j=1}^{l}\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}}.$
By the construction of the state $\Lambda^{s_{1},...,s_{l}}_{r}(\rho)$ we have
$c_{r}[\Lambda^{s_{1},...,s_{l}}_{r}(\rho)]_{A_{s}}\leq\rho_{A_{s}},\quad c_{r}=\mathrm{Tr}Q_{r}\rho,$
for any $s$. Hence, it follows from (29) that
$\sum_{s=1}^{m}\mathrm{Tr}G_{A_{s}}[\Lambda^{s_{1},...,s_{l}}_{r}(\rho)]_{A_{s}}\leq c^{-1}_{r}\sum_{s=1}^{m}\mathrm{Tr}G_{A_{s}}\rho_{A_{s}}=c^{-1}_{r}E_{S}\leq\frac{E_{S}}{1-\sum_{j=1}^{l}\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}}},$ (30)
where $E_{S}=E_{1}+...+E_{m}$.
By using Winter’s gentle measurement lemma (cf.[39]) we obtain
$\|\rho-\Lambda^{s_{1},...,s_{l}}_{r}(\rho)\|_{1}\leq 2\sqrt{\mathrm{Tr}\bar{Q}_{r}\rho}\leq 2\sqrt{\sum_{j=1}^{l}\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}}},$ (31)
where $\,\bar{Q}_{r}=I_{A_{1}...A_{n}}-Q_{r}\,$ and the last inequality follows from (29).
Let $\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ be a common faithful $I_{n}^{m}$-type continuity bound for the family $\mathcal{F}$. Let $f$ be any function in $\mathcal{F}$. Since $\,\sum_{s=1}^{m}\mathrm{Tr}G_{A_{s}}\rho_{A_{s}}=E_{S}$ and all the operators $G_{A_{1}}$,…,$G_{A_{m}}$ satisfy condition (8), by taking into account inequalities (30) and (31) we obtain
$|f(\rho)-f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))|\leq\widehat{B}_{\mathcal{F}}(\varepsilon_{r},2E_{S}|G_{\!A_{1}},..,G_{\!A_{m}}),\quad\varepsilon_{r}=\sqrt{\sum_{j=1}^{l}\mathrm{Tr}\bar{P}_{r}^{s_{j}}\rho_{A_{s_{j}}}},$ (32)
for all $r>r_{0}$, where $r_{0}$ is some natural number.
Since $\varepsilon_{r}\to+\infty$ as $r\to+\infty$, the r.h.s. of (32) tends to zero as $r\to+\infty$. By denoting the r.h.s. of (32) by $Y(r)$ for $r>r_{0}$ and assuming that
$Y(r)=\sup_{f\in\mathcal{F}}|f(\rho)-f(\Lambda^{s_{1},...,s_{l}}_{r}(\rho))|$ (33)
for $r\leq r_{0}$ we obtain the claim A of the theorem (by using the common continuity bound $\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ and inequality (30) it is easy to show that the supremum in (33) is finite for any $r$).
B) For each natural $s$ in $[1,m]$ let $\{g^{s}_{i}\}_{i=1}^{+\infty}$ be a sequence of nonnegative numbers such that
$\sum_{i=1}^{+\infty}\lambda^{s}_{i}g^{s}_{i}<+\infty\quad\textrm{and}\quad\lim_{\beta\to 0^{+}}\left[\sum_{i=1}^{+\infty}e^{-\beta g^{s}_{i}}\right]^{\beta}=1,$ (34)
where $\{\lambda^{s}_{i}\}_{i=1}^{+\infty}$ is the spectrum of the state $\rho_{A_{s}}$ taking in the non-increasing order. We may assume that $g_{1}^{s}=0$ for all $s$.
Then, by defining for each $s$ the positive operator $G_{\!A_{s}}$ on $\mathcal{H}_{A_{s}}$ by the formula (28) we may repeat all the arguments of the proof of claim A with the use of the common faithful $II_{n}^{m}$-type continuity bound $\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ for the family $\mathcal{F}$ at the last step (since now all the operators $G_{A_{1}}$,…,$G_{A_{m}}$ satisfy condition (10)).
By noting that the r.h.s. of (32) of tends to zero as $r\to+\infty$ and by constructing the sequence $Y(r)$ in the same way as in the proof of claim A we obtain claim B of the theorem.
The last claim of the theorem follows from the above proof. $\square$
Example 3. Let $f_{\Phi_{1}...\Phi_{n}}(\rho)=I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho)}$, where $\Phi_{1}:A_{1}\rightarrow B_{1}$,…, $\Phi_{n}:A_{n}\rightarrow B_{n}$ are arbitrary quantum channels. Inequality (10) in [31] (with trivial system $C$), upper bound (4), the nonnegativity of the quantum mutual information and its monotonicity under local channels imply that the function $f_{\Phi_{1}...\Phi_{n}}$ belongs to the class $L_{n}^{n-1}(2,n)$ for any channels $\Phi_{1},...,\Phi_{n}$. [^13] See the remark at the end of the Appendix. It follows that the quantity $\widehat{B}_{\mathcal{F}(2,n)}(\varepsilon,E|H_{A_{1}},..,H_{A_{n-1}})$ defined in (19) is a common faithful $II_{n}^{n-1}$-type continuity bound for the family =$\mathcal{F}=\{f_{\Phi_{1}...\Phi_{n}}\}_{\Phi_{1}...\Phi_{n}}$ (where $\Phi_{k}$ runs over the set of all channels from $A_{k}$ to $B_{k}$, $k=\overline{1,n}$). Thus, if the marginal states $\rho_{A_{1}}$,…,$\rho_{A_{n-1}}$ have the FA-property then, by Theorem 3B, for any subset $\{s_{1},...,s_{l}\}$ of $\{1,...,n\}$ there exists a vanishing sequence $\{Y(r)\}$ such that
$|I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho_{r})}-I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho)}|\leq Y(r)\qquad\forall r\in\mathbb{N},$ (35)
where $\rho_{r}=\Lambda^{s_{1},...,s_{l}}_{r}(\rho)$, for any channels $\Phi_{1},...,\Phi_{n}$.
The last assertion of Theorem 3 shows that inequality (35) remains valid with $\rho$ and $\rho_{r}=\Lambda^{s_{1},...,s_{l}}_{r}(\rho)$ replaced by $\sigma$ and $\sigma_{r}=\Lambda^{s_{1},...,s_{l}}_{r}(\sigma)$, where $\sigma$ is any state in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ such that $\sigma_{A_{k}}=\rho_{A_{k}}$ for each $k\in\{s_{1},...,s_{l}\}\cup\{1,...,n-1\}$.
By using expression (23), the Stinespring representation of the channels $\Phi_{1}$,…,$\Phi_{n-1}$ and Proposition 7 in [33, Section 4.2] it is easy to show the uniform continuity of the function
$(\Phi_{1},...,\Phi_{n})\mapsto I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho_{r})}$
on the set of all tuples $(\Phi_{1},...,\Phi_{n})$ of channels equipped with the topology of pairwise strong convergence, where $\rho_{r}=\Lambda^{1,...,n-1}_{r}(\rho)$, for any $r$. Thus, the above observation with $\{s_{1},...,s_{l}\}=\{1,...,n-1\}$ implies continuity of the function
$(\Phi_{1},...,\Phi_{n})\mapsto I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho)}$
on the set of all tuples $(\Phi_{1},...,\Phi_{n})$ of channels equipped with the topology of pairwise strong convergence provided that the FA-property holds for at least $n-1$ marginal states of the state $\rho$. By noting that the FA-property implies the finiteness of the entropy, this assertion can be also obtained from Proposition 3 in [30] proved in a completely different way.
4 On continuity of the regularized relative entropy of $\pi$-entanglement
In this section we apply Theorem 1 along with the results from [15] to obtain conditions of local continuity for the regularized relative entropy of $\pi$-entanglement of a state of $n$-partite system $A_{1}\ldots A_{n}$ corresponding to any given partition $\pi$ of $[n]\doteq\{1,\ldots,n\}$ (i.e. a finite collection $\{\pi(j)\}_{j\in J}$ of non-empty sets that do not intersect each other and together cover the whole set $[n]$). For a given $\pi$ we define the set of $\pi$-separable states by (cf.[25])
$\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})\doteq\mathrm{cl}\left(\mathrm{conv}\left(\left\{\bigotimes_{j\in J}|\psi_{j}\rangle\langle\psi_{j}|,\;|\psi_{j}\rangle\in\mathcal{H}_{A_{\pi(j)}},\;\langle\psi_{j}|\psi_{j}\rangle=1\right\}\right)\right),$
where $A_{\pi(j)}$ is the system obtained by joining those $A_{i}$ such that $i\in\pi(j)$, $\mathrm{conv}(\cdot)$ and $\mathrm{cl}(\cdot)$ denote, respectively, the convex hull and the closure w.r.t. the trace norm.
Let $\rho$ be a state of a $n$-partite quantum system $A_{1}\ldots A_{n}$. For any partition $\pi$, the relative entropy of $\pi$-entanglement of $\rho$ is defined as
$E_{R,\pi}(\rho)=\inf_{\sigma\in\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})}D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma).$ (36)
It is shown in [14] that this infimum is always attained and that the state at which it is attained is unique provided that $\rho$ is faithful state.
If $\pi$ is the finest partition $\left\{\{1\},\ldots,\{n\}\right\}$ then $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ coincides with the convex set $\mathfrak{S}_{\rm sep}(\mathcal{H}_{A_{1}...A_{n}})$ of all (fully) separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{m}})$ and hence $E_{R,\pi}$ is the relative entropy of entanglement $E_{R}$ in this case.
The relative entropy of $\pi$-entanglement possesses basic properties of entanglement measures (convexity, selective LOCC-monotonicity, asymptotic continuity, etc.) and satisfies the inequality
$pE_{R,\pi}(\rho)+(1-p)E_{R,\pi}(\sigma)\leq E_{R,\pi}(p\rho+(1-p)\sigma)+h_{2}(p),$ (37)
valid for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ and any $p\in[0,1]$, where $h_{2}$ is the binary entropy [41, 14]. It is bounded above by the sum of any $n-1$ marginal entropies, i.e.
$E_{R,\pi}(\rho)\leq\sum_{j=1}^{n-1}S(\rho_{A_{s_{j}}}),$ (38)
for any set $A_{s_{1}},...,A_{s_{n-1}}$ of $n-1$ subsystems [23],[14, the Appendix].
The inequalities (37) and (38) along with the convexity and nonnegativity of $E_{R,\pi}$ imply that $E_{R,\pi}\in L_{n}^{n-1}(1,1)$ and hence $E_{R,\pi}$ has a faithful $II_{n}^{n-1}$-type continuity bound due to the relation (18) in Section 2.3.
The relative entropy of $\pi$-entanglement is nonadditive on tensor products, its regularization is defined as
$E^{\infty}_{R,\pi}(\rho)=\lim_{k\rightarrow+\infty}k^{-1}E_{R,\pi}(\rho^{\otimes k})=\inf_{k\in\mathbb{N}}k^{-1}E_{R,\pi}(\rho^{\otimes k}),$ (39)
where $\rho^{\otimes k}$ is treated as a state of the $n$-partite quantum system $A^{k}_{1}...A^{k}_{n}$, where $A^{k}_{j}$ denotes the system obtained by joining $k$ copies of $A_{j}$. The above limit exists due to Fekete’s lemma (cf.[6]), because the relative entropy of $\pi$-entanglement is subadditive on tensor products.
It is shown in [14] that the relative entropy of $\pi$-entanglement $E_{R,\pi}$ is lower semicontinuous function on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{m}})$ for any partition $\pi$. Then a criterion of local continuity of $E_{R,\pi}$ is obtained in [15], which allows us to derive several verifiable conditions of local continuity of $E_{R,\pi}$ for any partition $\pi$, in particular, to prove that continuity of $E_{R,\pi}$ on a subset of $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ follows from continuity of the quantum mutual information on this subset (easy verifiable conditions for local continuity of the quantum mutual information are presented in [33, Section V]).
The regularized relative entropy of $\pi$-entanglement $E_{R,\pi}^{\infty}$ belongs to the class $II_{n}^{n-1}$ (in terms of Section 2.3) for any partition $\pi$: the faithful $II_{n}^{n-1}$-type continuity bound for $E_{R,\pi}^{\infty}$ is presented in Proposition 39 in [15]. This allows us to apply the approximation technique considered in Section 3 to the function $E^{\infty}_{R,\pi}$.
The following theorem contains results about analytical properties the regularized relative entropy of $\pi$-entanglement $E^{\infty}_{R,\pi}$ (proved by using Theorem 3B).
Theorem 2.  Let
$\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})=\left\{\hskip 1.0pt\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})\,|\,\rho_{A_{s}}\!\!\in\mathfrak{S}_{\rm\!\textsf{FA}\!}(\mathcal{H}_{A_{s}})\textrm{ for }\,n-1\,\textrm{ indexes}\hskip 1.0pts\hskip 1.0pt\right\}$ (40)
be the subset of $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of states $\rho$ such that the FA-property holds for at least $\hskip 1.0ptn-1$ marginal states of $\rho$. Let $\pi$ be any partition of $\{1,2,..,n\}$.
A) The function $E^{\infty}_{R,\pi}$ is finite and lower semicontinuous on the set $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. Moreover,
$\liminf_{k\to+\infty}E^{\infty}_{R}(\rho_{k})\geq E^{\infty}_{R}(\rho_{0})$ (41)
for arbitrary sequence $\{\rho_{k}\}\subset\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to a state $\rho_{0}\in\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$.
B) Let $\{\rho_{k}\}$ be an arbitrary sequence of states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to a state $\rho_{0}\in\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. The limit relation
$\lim_{k\to+\infty}E^{\infty}_{R,\pi}(\rho_{k})=E^{\infty}_{R,\pi}(\rho_{0})<+\infty$ (42)
holds provided one of the following conditions is valid: [^14] The finiteness of $\,E_{R,\pi}(\rho_{0})$ and $\,I(A_{1}\!:...:\!A_{n})_{\rho_{0}}$ follows from the condition $\rho_{0}\in\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$.
  1. $\rm a)$
    there exists a sequence $\{\omega_{k}\}\subset\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{m}})$ converging to a state $\omega_{0}\in\mathfrak{S}_{\rm sep}^{\pi}$ such that
    $\lim_{k\to+\infty}D(\rho_{k}\|\hskip 1.0pt\omega_{k})=D(\rho_{0}\|\hskip 1.0pt\omega_{0})<+\infty;$ (43)
  2. $\rm b)$
    there exists
    $\lim_{k\to+\infty}E_{R,\pi}(\rho_{k})=E_{R,\pi}(\rho_{0})<+\infty;$ (44)
  3. $\rm c)$
    there exists
    $\lim_{k\to+\infty}I(A_{1}\!:...:\!A_{n})_{\rho_{k}}=I(A_{1}\!:...:\!A_{n})_{\rho_{0}}<+\infty.$ (45)
Remark 6. The conditions $\rm a)$,$\,\rm b)$ and $\rm c)$ in Theorem 4 are related as follows: Theorem 1 in [15] states that $\,\rm a)\Rightarrow b)$ in general and that $\,\rm a)\Leftrightarrow b)$ if the state $\rho_{0}$ is faithful; Proposition 2 in [15] states that $\,\rm c)\Rightarrow b)$.
A criterion and different conditions for the validity of $\,\rm b)$ are described in [15].
Easy verifiable conditions for the validity of $\,\rm c)$ are presented in [30, Section 5].
Proof. A) Let $\rho_{0}$ be a state in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. Assume that the FA-property holds for the states $[\rho_{0}]_{A_{1}}$,…,$[\rho_{0}]_{A_{n-1}}$. The finiteness of $E^{\infty}_{R,\pi}(\rho_{0})$ follows from inequality (38), since the FA-property of a state implies finiteness of its entropy [32].
To prove relation (41) for $E^{\infty}_{R,\pi}$ we apply Theorem 3B with the family of maps $\{\widetilde{\Lambda}^{1,...,n-1}_{r}\}_{r}$ defined in (27). It is mentioned before that the function $E^{\infty}_{R,\pi}$ belongs to the class $II_{n}^{n-1}$. So, Theorem 3B along with Remark 3 imply that
$\lim_{r\to+\infty}E^{\infty}_{R,\pi}(\widetilde{\Lambda}^{1,...,n-1}_{r}(\rho_{0}))=E^{\infty}_{R,\pi}(\rho_{0}).$ (46)
Consider the function
$f_{r}(\rho)=E^{\infty}_{R,\pi}(\widetilde{\Lambda}^{1,...,n-1}_{r}(\rho))=E^{\infty}_{R,\pi}(\Phi^{1}_{r}\otimes...\otimes\Phi^{n-1}_{r}\otimes\mathrm{Id}_{A_{n}}(\rho)),$
on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, where $\Phi^{s}_{r}$, $s=1,2,...,n-1$, are the channels defined after (27). Since all the output states of the channel $\Phi^{s}_{r}$ are supported by the finite-dimensional subspace $P^{s}_{r}(\mathcal{H}_{A_{s}})$, the function $f_{r}$ is continuous on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ for each $r$ (this follows from the continuity bound (151) for $E^{\infty}_{R,\pi}$ in [14, Section V] obtained by generalizing the Winter arguments from [41]). Hence, $f_{*}=\sup_{r}f_{r}$ is a lower semicontinuous function on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$.
Since $f_{*}$ does not exceed the function $E^{\infty}_{R,\pi}$ by the monotonicity of $E^{\infty}_{R,\pi}$ under local channels, we have
$f_{*}(\rho_{k})\leq E^{\infty}_{R,\pi}(\rho_{k})\quad\forall k\qquad\textrm{and}\qquad f_{*}(\rho_{0})=E^{\infty}_{R,\pi}(\rho_{0}),$
where the equality follows from (46). Thus, the lower semicontinuity of $f_{*}$ implies (41).
B) By Remark 4 it suffices to prove that condition $\rm b)$ implies (42). Let $p\in\mathbb{N}$ be arbitrary. Since the cone generated by the set $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{p}_{1}...A^{p}_{m}})$ is a weak* closed subset of $\mathfrak{T}_{+}(\mathcal{H}_{A^{p}_{1}...A^{p}_{m}})$ for any partition $\pi$ [14] and the state $\omega^{\prime}\otimes\omega^{\prime\prime}$ lies in $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{p}_{1}...A^{p}_{m}})$ for any states $\omega^{\prime}\in\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{t}_{1}...A^{t}_{m}})$ and $\omega^{\prime\prime}\in\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{p-t}_{1}...A^{p-t}_{m}})$, $t=1,2,...,p-1$, the condition $\rm b)$ and the repeated use Corollary 2G in [15] allow us to show that
$\lim_{k\to+\infty}E_{R,\pi}(\rho^{\otimes p}_{k})=E_{R,\pi}(\rho^{\otimes p}_{0})<+\infty\quad\forall p\in\mathbb{N}.$
It follows that
$\limsup_{k\to+\infty}E_{R,\pi}^{\infty}(\rho_{k})\leq E_{R,\pi}^{\infty}(\rho_{0})<+\infty$ (47)
because $\,E_{R,\pi}^{\infty}(\rho_{k})=\inf_{p}p^{-1}E_{R,\pi}(\rho^{\otimes p}_{k})$ for all $k$. This relation and (41) imply (42). $\Box$
Remark 7. The proof of Theorem 4B shows that the limit relation (47) holds for an arbitrary sequence $\{\rho_{k}\}$ of states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to any state $\rho_{0}$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ provided that one of the conditions $\rm a)$,$\,\rm b)$ and $\rm c)$ in Theorem 4B is valid.
5 Other applications
In this section we consider several results about properties of the relative entropy of $\pi$-entanglement and its regularization obtained by using Theorem 3B in Section 3.
5.1 Simplified definition of $E_{R,\pi}$
The first result concerns the possibility to take the infimum in the definition (36) of the relative entropy of $\pi$-entanglement over a subset of $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of ”simple” $\pi$-separable states.
Proposition 3.  Let $A_{1},..,A_{n}$ be arbitrary quantum systems and $\,\pi=\{\pi(j)\}_{j=1}^{\ell}$ a given partition of the set $\{1,2,...,n\}$.
If $\rho$ is a state from the set $\,\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ defined in (40) then the infimum in the definition (36) of $E_{R,\pi}(\rho)$ can be taken only over all finitely-decomposable $\pi$-separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, i.e. states having the form
$\sigma=\sum_{k=1}^{m}p_{k}\hskip 1.0pt|\psi^{k}_{1}\rangle\langle\psi^{k}_{1}|\otimes...\otimes|\psi^{k}_{\ell}\rangle\langle\psi^{k}_{\ell}|,\quad\psi^{k}_{j}\in\mathcal{H}_{A_{\pi}(j)},\;\;\langle\psi^{k}_{j}|\psi^{k}_{j}\rangle=1,\;\;m<+\infty,$ (48)
where $\{p_{k}\}_{k=1}^{m}$ is a probability distribution and $A_{\pi(j)}$ is the system obtained by joining those $A_{i}$ such that $i\in\pi(j)$.
Proposition 5.1 implies, in particular, that we may define $E_{R,\pi}(\rho)$ for any state $\rho$ in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ by ignoring the existence of countably-non-decomposable $\pi$-separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, i.e. the states that can not represented as a countable convex mixture of pure $\pi$-separable states. This claim is not trivial, since in general there are no reasons to assume that the infimum in definition (36) of $E_{R,\pi}(\rho)$ can be taken over the dense subset of $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of countably decomposable $\pi$-separable states.
Proof. Let $\rho$ be a state in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. For each natural $s$ in $[1,n]$ let $P_{r}^{s}$ be the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues (taking the multiplicity into account). It is mentioned in Section 4 that the function $E_{R,\pi}$ belongs to the class $II_{n}^{n-1}$. Hence Theorem 3B implies that
$\lim_{r\to+\infty}E_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho),$ (49)
where $\rho_{r}=c_{r}^{-1}Q_{r}\rho\hskip 1.0ptQ_{r}$, $\,Q_{r}=P_{r}^{1}\otimes...\otimes P_{r}^{n}$, $c_{r}=\mathrm{Tr}Q_{r}\rho$.
Consider the function
$\widetilde{E}_{R,\pi}(\rho)=\inf_{\sigma\in\widetilde{\mathfrak{S}}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})}D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma),$
where $\widetilde{\mathfrak{S}}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$ is the set of finitely-decomposable separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$. Since the set $\widetilde{\mathfrak{S}}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$ is convex, the joint convexity of the relative entropy and Proposition 5.24 in [19] imply that the function $\widetilde{E}_{R,\pi}$ is convex and that the inequality (37) with $E_{R,\pi}$ replaced by $\widetilde{E}_{R,\pi}$ holds for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$. Note also that
$\widetilde{E}_{R,\pi}(\rho)\leq D(\rho\hskip 1.0pt\|\hskip 1.0pt\rho_{A_{1}}\otimes...\otimes\rho_{A_{n}})=I(A_{1}:...:A_{n})_{\rho}\leq 2\sum_{s=1}^{n-1}S(\rho_{A_{s}}),$
where the second inequality follows from (4). It follows that the function $\widetilde{E}_{R,\pi}$ belongs to the class $L_{n}^{n-1}(2,1)\subset II_{n}^{n-1}$ (see Section 2.3). Hence for any state $\rho$ in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ Theorem 3B implies that
$\lim_{r\to+\infty}\widetilde{E}_{R,\pi}(\rho_{r})=\widetilde{E}_{R,\pi}(\rho).$
By Lemma 5.1 below, we may assume in the definition of $E_{R,\pi}(\rho_{r})$ that $\rho_{r}$ is a state of a finite-dimensional $n$-partite system. Hence, $\widetilde{E}_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho_{r})$ for any $r$. So, the last limit relation and (49) show that $\widetilde{E}_{R,\pi}(\rho)=E_{R,\pi}(\rho)$ for any state $\rho$ in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$. $\Box$
Lemma 3. For an arbitrary state $\rho$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ the infimum in (36) can be taken over the set of all $\pi$-separable states in $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ supported by the subspace $\mathcal{H}^{1}_{\rho}\otimes...\otimes\mathcal{H}^{n}_{\rho}$, where $\mathcal{H}^{s}_{\rho}=\mathrm{supp}\rho_{A_{s}}$, $s=\overline{1,n}$.
Proof. Consider the channel
$\Phi(\varrho)=Q\varrho\hskip 1.0ptQ+[\mathrm{Tr}(I_{A_{1}...A_{n}}-Q)\varrho]\tau,\quad Q=P_{1}\otimes...\otimes P_{n},$
where $P_{s}$ is the projector on the subspace $\mathcal{H}^{s}_{\rho}$, $s=\overline{1,n}$, and $\tau$ is any $\pi$-separable state in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ supported by the subspace $\mathcal{H}^{1}_{\rho}\otimes...\otimes\mathcal{H}^{n}_{\rho}$. By monotonicity of the relative entropy we have
[…]
for any $\pi$-separable state $\sigma$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$. Since $\Phi(\sigma)$ is a $\pi$-separable state supported by the subspace $\mathcal{H}^{1}_{\rho}\otimes...\otimes\mathcal{H}^{n}_{\rho}$, the above inequality implies the assertion of the lemma. $\square$
Corollary 1. For an arbitrary $\pi$-separable state $\rho$ in $\,\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ there is a sequence $\{\rho_{k}\}$ of $\pi$-separable states having the form (48) such that
$\lim_{k\to+\infty}D(\rho\hskip 1.0pt\|\rho_{k})=0.$
5.2 Finite-dimensional approximation of $E_{R}$ and $E_{R}^{\infty}$ and its use
Many results describing properties of the relative entropy of $\pi$-entanglement and its regularization in finite-dimensional multipartite quantum systems remain valid in the infinite-dimensional case under some additional conditions. A simple way to establish the validity of these results is to use the approximation technique based on Theorem 4 and the results in [15].
We illustrate this approach restricting attention to the relative entropy of entanglement $E_{R}$ and its regularization $E_{R}^{\infty}$ defined by formulae (36) and (39) in which $\pi$ is the finest partition $\left\{\{1\},\ldots,\{n\}\right\}$ (in this case $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ coincides with the convex set $\mathfrak{S}_{\rm sep}(\mathcal{H}_{A_{1}...A_{n}})$ of all fully separable states).
Proposition 4.  Let $\rho$ be a state from the set $\,\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ defined in (40). Let $\rho_{k}=Q_{k}\rho\hskip 1.0ptQ_{k}[\mathrm{Tr}Q_{k}\rho\hskip 1.0pt]^{-1}$, $Q_{k}=P_{k}^{1}\otimes...\otimes P_{k}^{n}$, where $\{P_{k}^{1}\}\subset\mathfrak{B}(\mathcal{H}_{A_{1}})$,…, $\{P_{k}^{n}\}\subset\mathfrak{B}(\mathcal{H}_{A_{n}})$ are arbitrary sequences of projectors strongly converging to the unit operators $I_{A_{1}}$,…,$I_{A_{n}}$ correspondingly. Then
$E^{*}_{R}(\rho_{A_{1}...A_{m}})=\lim_{k\to+\infty}E^{*}_{R}([\rho_{k}]_{A_{1}...A_{m}}),\quad E_{R}^{*}=E_{R},E_{R}^{\infty},$ (50)
for any $\,m=2,3,...,n$, where $E_{R}$ and $E_{R}^{\infty}$ denote, respectively, the relative entropy of entanglement and its regularization of a state of the system $A_{1}...A_{m}$.
To guarantee the validity of (50) with $E_{R}^{*}=E_{R}$ it suffices to require that $\rho$ is any state in $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ having at least $\,n-1$ marginal states with finite entropy.
Proof. Assume that $\rho$ is any state in $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ having at least $\,n-1$ marginal states with finite entropy. This assumptions holds for any state in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$.
By using the monotonicity of the quantum mutual information (QMI) under local quantum operations it is easy to show that
$c_{k}I(A_{1}\!:...:\!A_{n})_{\rho_{k}}\leq I(A_{1}\!:...:\!A_{n})_{\rho},\quad c_{k}=\mathrm{Tr}Q_{k}\rho.$
Since $c_{k}\to 0$ as $k\to+\infty$, this inequality and the lower semicontinuity of QMI imply that
$\lim_{k\to+\infty}I(A_{1}\!:...:\!A_{n})_{\rho_{k}}=I(A_{1}\!:...:\!A_{n})_{\rho}<+\infty.$ (51)
The finiteness of $I(A_{1}\!:...:\!A_{n})_{\rho}$ follows from the upper bound (4).
By Proposition 3 in [30] the limit relation (52) shows that
$\lim_{k\to+\infty}I(A_{1}\!:...:\!A_{m})_{\rho_{k}}=I(A_{1}\!:...:\!A_{m})_{\rho}<+\infty,\quad m=2,3,..$ (52)
If $E_{R}^{*}=E_{R}$ then (50) follows from (52) by Proposition 2 in [15]. If $E_{R}^{*}=E_{R}^{\infty}$ and $\rho$ is a state in $\,\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ then (50) follows from (51) by Theorem 4B in Section 4. $\Box$
Below we consider examples of using Proposition 5.2.
Example 4. Lemma 5 in [21] and the additivity of the entropy imply that
$E_{R}^{\infty}(\rho)\geq-S(A_{i}|A_{j})_{\rho},\quad(i,j)=(1,2),(2,1),$ (53)
for any state $\rho$ of a bipartite finite-dimensional quantum system $A_{1}A_{2}$.
Proposition 5.2 allows us to show that the inequalities in (53) remain valid for any state $\rho$ of a bipartite infinite-dimensional system provided that the state $\rho_{A_{i}}$ has theFA-property [^15] This assumption implies, by Theorem 1 in [32], the finiteness of the extended conditional entropy $S(A_{i}|A_{j})_{\rho}$ defined in (3). and $S(\cdot|\cdot)$ is the extended conditional entropy defined in (3). Indeed, it suffices to show that
$S(A_{1}|A_{2})_{\rho}=\lim_{k\to+\infty}S(A_{1}|A_{2})_{\rho_{k}},\quad\rho_{k}=Q_{k}\rho\hskip 1.0ptQ_{k}[\mathrm{Tr}Q_{k}\rho\hskip 1.0pt]^{-1},\quad Q_{k}=P_{k}^{1}\otimes P_{k}^{2},$
where $P_{k}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues, $s=1,2$. This can be done by using Theorem 3B, since the function $\rho\mapsto S(A_{1}|A_{2})_{\rho}$ belongs to the class $L_{2}^{1}(2,1)\subset II_{2}^{1}$ (see Section 2.3).
Note that direct proof of (53) in the infinite-dimensional case requires technical efforts (especially, if $S(\rho)=S(\rho_{A_{j}})=+\infty$).
Example 5. It is shown in [23] that
$E_{R}^{*}(\rho)\geq E_{R}^{*}(\rho_{A_{i}A_{j}})+S(\rho_{A_{i}A_{j}}),\;\;E_{R}^{*}=E_{R},E_{R}^{\infty},\;\;(i,j)=(1,2),(2,3),(3,1),$ (54)
for any pure state $\rho$ in a tripartite finite-dimensional quantum system $A_{1}A_{2}A_{3}$.
Proposition 5.2 allows us to show that all the inequalities in (54) with $E_{R}^{*}=E_{R}$ (resp. with $E_{R}^{*}=E_{R}^{\infty}$) remain valid for any pure state $\rho$ of a tripartite infinite-dimensional system provided that any two of the states $\rho_{A_{1}}$, $\rho_{A_{2}}$ and $\rho_{A_{3}}$ have finite entropy (resp. have the FA-property). [^16] This assumption and the purity of the state $\rho$ imply, by Theorem 1 in [32], the finiteness of all the entropies $S(\rho_{A_{i}A_{j}})$ involved in (54). Indeed, it suffices to show that
$S(\rho_{A_{1}A_{2}})=\lim_{k\to+\infty}S([\rho_{k}]_{A_{1}A_{2}}),\quad\rho_{k}=Q_{k}\rho\hskip 1.0ptQ_{k}[\mathrm{Tr}Q_{k}\rho\hskip 1.0pt]^{-1},\quad Q_{k}=P_{k}^{1}\otimes P_{k}^{2}\otimes P_{k}^{3},$
where $P_{k}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues, $s=1,2,3$. This can be done by noting that $S(\rho_{A_{1}A_{2}})=S(\rho_{A_{3}})$ and $S([\rho_{k}]_{A_{1}A_{2}})=S([\rho_{k}]_{A_{3}})$ as the states $\rho$ and $\rho_{k}$ are pure. Since $c_{k}[\rho_{k}]_{A_{3}}\leq\rho_{A_{3}}$ for each $k$ for some positive number $c_{k}$ tending to $1$ as $k\to+\infty$, the required limit relation can be proved easily by using concavity and lower semicontinuity of the entropy.
5.3 On the energy-constrained versions of $E_{R,\pi}$
Dealing with infinite-dimensional quantum systems we have to take into account the existence of quantum states with infinite energy which can not be produced in a physical experiment. This motivates an idea to define the relative entropy of $\pi$-entanglement of a state $\rho$ with finite energy by formula (36) in which the infimum is taken over all $\pi$-separable states with the energy not exceeding some (sufficiently large) bound $E$ [4]. This gives the following energy-constrained version of the relative entropy of $\pi$-entanglement
$E^{H_{\!A^{n}}}_{R,\pi}(\rho\hskip 1.0pt|E)=\inf_{\sigma\in\mathfrak{S}_{\mathrm{sep}}^{\pi}(\mathcal{H}_{A_{1}...A_{n}}),\mathrm{Tr}H_{\!A^{n}}\sigma\leq E}D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma),\quad$ (55)
where $H_{\!A^{n}}$ is the Hamiltonian of a $n$-partite system $A^{n}\doteq A_{1}...A_{n}$ [4]. [^17] In [4] the case when $E_{R,\pi}=E_{R}$ (i.e. when $\pi$ is the finest partition of $\{1,2,...,n\}$) is considered.
An obvious drawback of definition (55) is its dependance of the bound $E$. Note also that $E^{H_{\!A^{n}}}_{R,\pi}(\rho\hskip 1.0pt|E)\neq 0$ for any $\pi$-separable state $\rho$ such that $\mathrm{Tr}H_{\!A^{n}}\rho>E$ and that the monotonicity of $E^{H_{\!A^{n}}}_{R,\pi}(\rho\hskip 1.0pt|E)$ under local operations can be shown only under the appropriate restrictions on the energy amplification factors of such operations.
A less restrictive way to take the energy constraints into account is to take the infimum in formula (36) over all $\pi$-separable states with finite energy, i.e. to consider the following quantity
$E^{H_{\!A^{n}}}_{R,\pi}(\rho)=\inf_{\sigma\in\mathfrak{S}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}}),\mathrm{Tr}H_{\!A^{n}}\sigma<+\infty}D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma)=\lim_{E\rightarrow+\infty}E^{H_{\!A^{n}}}_{R,\pi}(\rho\hskip 1.0pt|E),$ (56)
where $H_{\!A^{n}}$ is the Hamiltonian of $A^{n}\doteq A_{1}...A_{n}$. The limit in (56) can be replaced by the infimum over all $E>0$, since the function $E\mapsto E^{H_{\!A^{n}}}_{R,\pi}(\rho\hskip 1.0pt|E)$ is non-increasing. If the Hamiltonian $H_{\!A^{n}}$ has the form
$H_{\!A^{n}}=H_{A_{1}}\otimes I_{A_{2}}\otimes...\otimes I_{A_{n}}+\cdots+I_{A_{1}}\otimes...\otimes I_{A_{n-1}}\otimes H_{A_{n}},$ (57)
where $H_{A_{s}}$ is a densely defined positive operator on $\mathcal{H}_{A_{s}}$, $s=\overline{1,n}$, then the set of $\pi$-separable states with finite energy is dense in $\mathfrak{S}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$. So, in this case it is reasonable to assume the coincidence of $E^{H_{\!A^{n}}}_{R,\pi}(\rho)$ and $E_{R,\pi}(\rho)$ for any state $\rho$ with finite energy. [^18] This coincidence is not obvious, since the infima of a lower semicontinuous function over a closed set and over a dense subset if this set may be different. In the following proposition we establish the validity of this assumption under the particular condition on the Hamiltonians $H_{A_{1}}$,…,$H_{A_{n}}$.
Proposition 5.  Let $H_{\!A^{n}}$ be the Hamiltonian of a composite system $A_{1}...A_{n}$ expressed by formula (57) via the Hamiltonians $H_{A_{1}}$,…,$H_{A_{n}}$ of the subsystems $A_{1}$,…,$A_{n}$. If at least $\,n-1\,$ of them satisfy condition (10) then
$E^{H_{\!A^{n}}}_{R,\pi}(\rho)=E_{R,\pi}(\rho)$ (58)
for any state $\rho$ such that $\mathrm{Tr}H_{\!A^{n}}\rho=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\rho_{A_{s}}<+\infty$. Moreover, for any such state $\rho$ the infimum in the definition (36) of $E_{R,\pi}(\rho)$ can be taken only over all finitely-decomposable separable states $\sigma$ (i.e. the states having the form (48)) such that $\mathrm{Tr}H_{\!A^{n}}\sigma=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\sigma_{A_{s}}<+\infty$.
Proof. Assume that the Hamiltonians $H_{A_{1}}$,…,$H_{A_{n-1}}$ satisfy condition (10). Let $\mathfrak{S}_{0}$ be the convex subset of $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of all states $\rho$ with finite value of $\mathrm{Tr}H_{\!A^{n}}\rho=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\rho$. Consider the function
$\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho)=\inf_{\sigma\in\mathfrak{S}^{\mathrm{\pi,f}}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})\cap\mathfrak{S}_{0}}D(\rho\hskip 1.0pt\|\hskip 1.0pt\sigma),$
on the set $\mathfrak{S}_{0}$, where $\mathfrak{S}^{\mathrm{\pi,f}}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$ is the set of finitely-decomposable $\pi$-separable states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$. Since the set $\mathfrak{S}^{\mathrm{\pi,f}}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})\cap\mathfrak{S}_{0}$ is convex, the joint convexity of the relative entropy and Proposition 5.24 in [19] imply that the function $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}$ is convex and that the inequality (37) with $E_{R,\pi}$ replaced by $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}$ holds for any states $\rho$ and $\sigma$ in $\mathfrak{S}_{0}$. Since
$\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho)\leq D(\rho\hskip 1.0pt\|\hskip 1.0pt\rho_{A_{1}}\otimes...\otimes\rho_{A_{n}})=I(A_{1}:...:A_{n})_{\rho}\leq 2\sum_{s=1}^{n-1}S(\rho_{A_{s}})\quad\forall\rho\in\mathfrak{S}_{0},$
where the second inequality follows from (4), the function $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}$ belongs to the class $L_{n}^{n-1}(2,1|\hskip 1.0pt\mathfrak{S}_{0})$ in terms of Remark 9 in [31]. It is clear that $\,E_{R,\pi}(\rho)\leq E^{H_{\!A^{n}}}_{R,\pi}(\rho)\leq\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho)\,$ for any state $\rho$ in $\mathfrak{S}_{0}$.
For each natural $s$ in $[1,n-1]$ let $P_{r}^{s}$ be the spectral projector of the operator $H_{A_{s}}$ corresponding to its $r$ minimal eigenvalues (taking the multiplicity into account). Let $P_{r}^{n}=\sum_{i=1}^{r}|i\rangle\langle i|$, where $\{|i\rangle\}$ is a basis in $\mathcal{H}_{A_{n}}$ such that $\langle i|H_{A_{n}}|i\rangle<+\infty$ for all $i$.
By the proof of Theorem 1 in [28] the set $\mathfrak{S}_{0}$ has the invariance property stated in Remark 9 in [31]. By this remark Theorem 5 in [31] is generalized to all the functions from the class $L_{n}^{n-1}(2,1|\hskip 1.0pt\mathfrak{S}_{0})$ containing the function $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}$. Since the function $E_{R,\pi}$ belongs to the class $L_{n}^{n-1}(1,1)$, Theorem 5 in [31], its generalization mentioned before and the arguments from the proof of Theorem 3 in Section 3 show that
$\lim_{r\to+\infty}E_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho)\quad\textrm{and}\quad\lim_{r\to+\infty}\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho_{r})=\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho),\quad\forall\rho\in\mathfrak{S}_{0},$ (59)
where $\rho_{r}=c_{r}^{-1}Q_{r}\rho\hskip 1.0ptQ_{r}$, $\,Q_{r}=P_{r}^{1}\otimes...\otimes P_{r}^{n}$, $c_{r}=\mathrm{Tr}Q_{r}\rho$. Hence, to prove that $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho)=E^{H_{\!A^{n}}}_{R,\pi}(\rho)=E_{R,\pi}(\rho)$ it suffices to show that $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho_{r})$ for any $r$. This can be done by using Lemma 5.1 in Section 5.1 and by noting that $\mathrm{Tr}H_{\!A^{n}}\sigma=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\sigma_{A_{s}}<+\infty$ for any $\pi$-separable state $\sigma$ supported by the subspace $P_{r}^{1}\otimes...\otimes P_{r}^{n}(\mathcal{H}_{A_{1}...A_{n}})$. $\square$
Appendix: On the classes $\widehat{L}^{m}_{n}(C,D)$ and $N^{m}_{n,s}(C,D)$
Below a description of the classes $\widehat{L}^{m}_{n}(C,D)$ and $N^{m}_{n,s}(C,D)$ of functions on the set of states of a composite quantum system $A_{1}...A_{n}$ is presented along with a list of basic characteristics belonging to these classes. The functions in $\widehat{L}^{m}_{n}(C,D)\cup N^{m}_{n,s}(C,D)$ has a faithful common $II^{m}_{n}$-type continuity bound under the energy-type constraints in the subsystems $A_{1}$,..,$A_{m}$ given by expression (19). [^19] If $A_{1}$,…,$A_{m}$ are finite-dimensional quantum systems then a faithful common Fannes’ type continuity bound for the functions in $\widehat{L}^{m}_{n}(C,D)\cup N^{m}_{n,s}(C,D)$ is given by Theorem 3 in [33, Section 3].
Let $m\leq n$ and $L^{m}_{n}(C,D)$ be the class of all (finite) functions $f$ on the set [^20] We assume that $A_{1}$,…,$A_{n}$ are infinite-dimensional quantum systems.
$\mathfrak{S}_{m}(\mathcal{H}_{A_{1}..A_{n}})\doteq\left\{\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})\,|\,S(\rho_{A_{1}}),...,S(\rho_{A_{m}})<+\infty\hskip 1.0pt\right\}$
such that
$-c^{-}_{f}C_{m}(\rho)\leq f(\rho)\leq c^{+}_{f}C_{m}(\rho),\quad\quad C_{m}(\rho)=\sum_{s=1}^{m}S(\rho_{A_{s}}),$ (60)
for any state $\rho$ in $\mathfrak{S}_{m}(\mathcal{H}_{A_{1}..A_{n}})$ and
$-d^{-}_{f}h_{2}(p)\leq f(p\rho+(1-p)\sigma)-pf(\rho)-(1-p)f(\sigma)\leq d^{+}_{f}h_{2}(p)$ (61)
for any states $\rho$ and $\sigma$ in $\mathfrak{S}_{m}(\mathcal{H}_{A_{1}..A_{n}})$ and any $p\in(0,1)$, where $h_{2}$ is the binary entropy (defined after (2)), $c^{\pm}_{f}$ and $d^{\pm}_{f}$ are nonnegative numbers such that $c^{-}_{f}+c^{+}_{f}=C$ and $d^{-}_{f}+d^{+}_{f}=D$.
Let $\widehat{L}^{m}_{n}(C,D)$ be the class containing all functions in $L^{m}_{n}(C,D)$ and all the functions of one of the forms
$f(\rho)=\inf_{\lambda}f_{\lambda}(\rho)\quad\textrm{and}\quad f(\rho)=\sup_{\lambda}f_{\lambda}(\rho),$
where $\{f_{\lambda}\}$ is a family of functions in $L^{m}_{n}(C,D)$.
Let $N^{m}_{n,s}(C,D)$, $s=1,2,3$, be the class of all functions $f$ on the set $\mathfrak{S}_{m}(\mathcal{H}_{A_{1}..A_{n}})$ defined by one of the expressions
$f(\rho)\doteq\inf_{\hat{\rho}\in\mathfrak{M}_{s}(\rho)}h(\hat{\rho})\quad\textrm{and}\quad f(\rho)\doteq\sup_{\hat{\rho}\in\mathfrak{M}_{s}(\rho)}h(\hat{\rho})$
via a particular function $h$ in $\widehat{L}^{m}_{n+l}(C,D)$ for some $l>0$, where:
  • $\mathfrak{M}_{1}(\rho)$ is the set of all extensions of $\rho$ to a state of $A_{1}...A_{n+l}$;
  • $\mathfrak{M}_{2}(\rho)$ is the set of all extensions of $\rho$ having the form
    $\hat{\rho}=\sum_{i}p_{i}\rho_{i}\otimes|i\rangle\langle i|,$ (62)
    where $\{\rho_{i}\}$ is a collection of states in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$, $\{p_{i}\}$ is a probability distribution and $\{|i\rangle\}$ is an orthonormal basis in $\mathcal{H}_{A_{n+1}}$ (in this case $l=1$);
  • $\mathfrak{M}_{3}(\rho)$ is the set of all extensions of $\rho$ having the form (62) in which $\{\rho_{i}\}$ is a collection of pure states in $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$.
Below is a (incomplete) list of characteristics of a quantum system $A_{1}...A_{n}$, indicating the classes to which they belong with reference to the justification:
  • the von Neumann entropy $S(\rho)$ of a state $\rho$ of $A_{1}$ ($L^{1}_{1}(1,1)$, the concavity of $S$ and inequality (2));
  • the (extended) quantum conditional entropy $S(A_{1}|A_{2})$ defined in (3) ($L^{1}_{2}(2,1)$, [33, Section 4.1]);
  • the quantum mutual information $I(A_{1}\!:...:\!A_{n})$ ($L^{n}_{n}(1,n)$ and $L^{n-1}_{n}(2,n)$, [33, Section 4.2]);
  • the (extended) quantum conditional mutual information $I(A_{1}\!:...:\!A_{n}|A_{n+1})$ ($L^{n-1}_{n+1}(2,n)$ and $L^{n}_{n+1}(2-1/n,n)$, [33, Section 4.2]);
  • the one-way classical correlation in $A_{1}A_{2}$ with measured system $A_{1}$ ($\widehat{L}^{1}_{2}(1,2)$, [33, Section 4.3]);
  • the one-way classical correlation in $A_{1}A_{2}$ with measured system $A_{2}$ ($\widehat{L}^{1}_{2}(1,2)$, [33, Section 4.3]);
  • the quantum discord in $A_{1}A_{2}$ with measured system $A_{1}$ ($\widehat{L}^{1}_{2}(1,2)$, [33, Section 4.3]);
  • the quantum discord in $A_{1}A_{2}$ with measured system $A_{2}$ ($\widehat{L}^{1}_{2}(2,2)$, [33, Section 4.3]);
  • the constrained Holevo capacity of the partial trace over $A_{2}$ in $A_{1}A_{2}$ ($N^{1}_{2,2}(1,2)$, [33, Section 4.5]);
  • the constrained Holevo capacity $\bar{C}(\Phi,\rho)$ of a quantum channel $\Phi:A_{1}\rightarrow B$ at a state $\rho$ of $A_{1}$ ($N^{1}_{1,2}(1,2)$, [33, Section 4.5]); [^21] This follows, due to relation (74) in [34], from the fact that the constrained Holevo capacity of the partial trace over $E$ in a bipartite system $BE$ belongs to the class $N^{1}_{1,2}(1,2)$ in the setting $A_{1}=BE$ [33, Section 4.5].
  • the mutual information $I(\Phi,\rho)$ of a quantum channel $\Phi:A_{1}\rightarrow B$ at state $\rho$ of $A_{1}$ ($L^{1}_{1}(2,2)$, [29, Section 4.3]);
  • the coherent information $I_{c}(\Phi,\rho)$ of a quantum channel $\Phi:A_{1}\rightarrow B$ at state $\rho$ of $A_{1}$ ($L^{1}_{1}(2,2)$, [29, Section 4.3]);
  • the entropy reduction of a measurement in $A_{1}$ ($L^{1}_{1}(1,1)$, [34, Section III]); [^22] We mean the entropy reduction of a POVM (rather than the entropy reduction of an instrument) as it is defined, for example, in [34, Section III]. The arguments from the proof of Proposition 1 in [34] show the validity of inequality (61) for the entropy reduction with $d^{-}_{f}=0$ and $d^{+}_{f}=1$, the inequality (60) with $c^{-}_{f}=0$ and $c^{+}_{f}=1$ follows from the nonnegativity of the entropy reduction and its obvious upper bound.
  • the entropy reduction of a local measurement in $A_{1}A_{2}$ with measured system $A_{1}$ ($L^{1}_{2}(1,1)$, [34, Section III]);
  • the entanglement of formation in $A_{1}A_{2}$ ($N^{1}_{2,3}(1,1)$, [33, Section 4.4]);
  • the relative entropy of $\pi$-entanglement in $A_{1}...A_{n}$ for any set of partitions $\pi$ ($L^{n-1}_{n}(1,1)$ and $L^{n}_{n}(1-1/n,1)$, [14, Section 5]);
  • the relative entropy of NPT-entanglement in $A_{1}A_{2}$ ($L^{1}_{2}(1,1)$ and $L^{2}_{2}(1/2,1)$, [14, Section 5]);
  • the Rains bound in $A_{1}A_{2}$ ($L^{1}_{2}(1,1)$ and $L^{2}_{2}(1/2,1)$, [14, Section 5]);
  • the squashed entanglement in $A_{1}...A_{n}$ ($N^{n-1}_{n,1}(1,n/2)$ and $N^{n}_{n,1}(1-1/n,n/2)$, [33, Section 4.6]);
  • the c-squashed entanglement in $A_{1}...A_{n}$ ($N^{n-1}_{n,2}(1,n/2)$ and $N^{n}_{n,2}(1-1/n,n/2)$, [33, Section 4.6]);
  • the conditional entanglement of mutual information in $A_{1}...A_{n}$ ($N^{n}_{n,1}(1,(n+1)/2)$, [31, Section IV-C]).
Note also that for any nonnegative characteristic $f$ of a multipartite quantum system $A_{1}...A_{n}$ not increasing under local channels acting on the subsystems $A_{1}$,…,$A_{m}$ the implications
$f\in L^{m}_{n}(C,D)\;\Rightarrow\;f\circ\Lambda\in L^{m}_{n}(C,D)\quad\textrm{and}\quad f\in\widehat{L}^{m}_{n}(C,D)\;\Rightarrow\;f\circ\Lambda\in\widehat{L}^{m}_{n}(C,D)$
hold for any channel
$\Lambda=\Phi_{1}\otimes\Phi_{2}\otimes\cdots\otimes\Phi_{n},$ (63)
where $\Phi_{k}$ is a channel from the system $A_{k}$ to any system $B_{k}$, $k=\overline{1,n}$ [31, Section V]. [^23] We assume that the function $f$ is defined on the set of states of any $n$-partite system, in particular, the system $B_{1}...B_{n}$.
This implies, for example, that the family of functions $f_{\Lambda}(\rho)=I(B_{1}\!:\!B_{2}|B_{3})_{\Lambda(\rho)}$, where $\Lambda$ runs over the set of all channels of the form (63) with $n=3$, is contained in the class $L^{1}_{3}(2,2)$ (because the function $f(\rho)=I(A_{1}\!:\!A_{2}|A_{3})_{\rho}$ is nonnegative, belongs to the class $L^{1}_{3}(2,2)$ and does not increase under channels acting on $A_{1}$). Hence, this family has a common faithful $II^{1}_{3}$-type continuity bound (in terms of Section 2.3).
I am grateful to A.S.Holevo, G.G.Amosov and E.R.Loubenets for useful discussion. I am also grateful to L.Lami for the collaboration that motivated this research.